# 2Instituto de Física de Partículas y del Cosmos (IPARCOS), Universidad Complutense de Madrid, E-28040 Madrid, Spain

### v1

@@ Abstract
Methods. We analyze first 18 years of Fermi-LAT data using 30-day binning. To isolate variability on QPO timescales, we apply Singular Spectrum Analysis to mitigate red-noise effects and use a Moving Block Bootstrap approach to quantify correlations between photon flux and photon index while accounting for temporal autocorrelation.
@@ 1 Introduction
In this Letter, we investigate the physical link between the 1.7 yr QPO and the transient spectral hardening in PKS 2155-304, defining spectral hardening as a departure from a single power-law gamma-ray spectrum, manifested as high-energy upturns in the GeV band. Using 18 years of Fermi-LAT data, we characterize the flux vs. photon index behavior and map the hardening event onto the phase of the oscillation. We find that the bulk variability is atypical for an HSP, showing a softer-when-brighter trend with no significant correlation between photon index and phase, consistent with the coexistence of stochastic plasma-driven activity and a geometric modulation. In addition, the uncommon spectral hardening event is phase-locked to the QPO trough.
@@ 2 Data and Methodology
We analyzed the $\gamma$-ray emission of PKS 2155-304 using approximately 18 years of Fermi-LAT data (August 2008–January 2026; MJD 54698–61058) obtained from the Fermi-LAT Light Curve Repository [^1] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/LightCurveRepository/ (LCR; Abdollahi et al., 2023). To characterize variability on the timescales of the reported QPO, we adopted a 30-day binning with a minimum test statistic threshold of $TS\geq 4$ per bin and allowed the photon index ($\Gamma$) to vary freely in each bin.
@@ 3.1 Global Spectral Behavior
We first characterized the global spectral variability of PKS 2155-304 over the full 18-year baseline to establish a physical reference for the QPO modulation (Figure 1). Figure 2 illustrates the relationship between the photon flux and the photon index. In contrast to the harder-when-brighter trend typical of most HSP blazars, PKS 2155-304 shows a statistically significant positive correlation.
@@
This trend is consistent with a geometric origin, such as periodic variations in the Doppler factor. In this scenario, the flux maximum corresponds to the smallest viewing angle, where the Doppler boosting effectively amplifies the dominant, softer portion of the spectral energy distribution (such as the cooled electron population or the jet sheath) more strongly than the hard core, resulting in a softening of the observed index during peaks.
@@
Figure 2: Spectral variability of PKS 2155-304. Left: Photon index vs. photon flux, showing a positive correlation ($\rho\approx 0.3$), weak but robust, indicative of an atypical softer-when-brighter trend. Right: Phase-resolved spectral evolution folded at the 1.7 yr period, with the 18-year baseline variability (green dots) consistent with an achromatic trend given the lack of significant correlation between photon index and phase ($\rho=0.036$, $p=0.60$), and the two consecutive 30-day bins of the extreme spectral hardening event reported by (Dinesh et al., 2025, red dots).
@@ 3.2 Phase-Locked Spectral Hardening
To quantify the significance of this alignment, we performed a phase-scrambling Monte Carlo test ($10^{4}$ realizations) in which the MJD window of the hardening event was randomly reassigned to any valid 30-day bin within the 18-year baseline, and the resulting phase was recorded. The probability of a random placement falling within $|\Delta\phi|\leq 0.03$ of the trough center is $p\approx 0.059$. Combining this with the independent flux–index correlation result ($p\approx 1.8\times 10^{-5}$) under the assumption of statistical independence, the joint probability of both features arising by chance is $p_{\mathrm{joint}}\approx 1.1\times 10^{-6}$, strongly disfavoring a coincidental alignment. We interpret the physical origin of this phase-locked behavior in the following section. Note that the assumption of independence is reasonable here, as the global flux-index correlation reflects the overall chromatic behavior of the source across the full 18-year dataset, while the phase alignment test specifically evaluates the timing of a single transient hardening event relative to the periodic cycle. These probe distinct physical and statistical aspects of the variability, with no evident coupling in the underlying data generation process.
@@ 4.1 The Geometric Masking Scenario
The phase alignment of the spectral hardening event with the QPO trough ($\phi\approx 0.53$) suggests a conditional visibility mechanism. We propose that the observational rarity of spectral hardening in HSP blazars ($0.1\%$; Dinesh et al. 2025) is not necessarily an indication that extreme particle acceleration events are uncommon, but rather that they are frequently masked.
In a geometric QPO scenario, flux peaks are driven by maximum relativistic beaming. Our finding of a global softer-when-brighter trend ($\rho\approx 0.3$) indicates that these boosted states are dominated by a soft photon field, likely originating from cooled electron populations or the jet sheath. This may redefine the high-flux state not as a period of intrinsic quiescence, but as a regime of maximum geometric masking. The jet may host frequent plasma-driven particle acceleration events that remain observationally unresolved, overwhelmed by the intensified soft background. During these phases, the spectral signature of intrinsic acceleration, such as shocks or magnetic reconnection, is suppressed by the boosted soft emission. Conversely, at the QPO trough, the Doppler factor is minimized, effectively suppressing the geometric boost and reducing this soft background. This reduces the geometric masking, allowing hard-spectrum emission from transient jet instabilities to dominate the observed flux and making the hardening detectable.
@@
An important corollary of this scenario addresses the rarity of the detected event. While the masking effect is minimized periodically every 1.7 yr, Dinesh et al. (2025) reported only one significant spectral hardening episode over the 18-year baseline. This apparent scarcity is likely due to two factors. First, the geometric minimum is a necessary condition for visibility, not a driver of acceleration; an intrinsic stochastic shock must still coincide with the narrow temporal window of the trough. Second, the blind search strategy employed by Dinesh et al. (2025) relied on high-significance deviations which favor high-flux states. Since QPO troughs are by definition low-flux intervals, weaker or shorter-duration hardening events may have failed to meet the strict global detection thresholds. We predict that a targeted search specifically focused on the QPO troughs of PKS 2155-304 will reveal a population of sub-threshold hardening events that were missed by previous blind surveys.
@@ 4.2 Constraints on Jet Structure and Extreme States
The coexistence of a stable geometric modulation and transient spectral hardening provides a unique constraint on the jet’s spatial architecture. While purely plasma-driven QPO models, such as recurrent shocks (e.g., Kirk et al., 1998), could in principle produce similar periodicities, they generally predict harder-when-brighter spectral evolution, which is inconsistent with our observed softer-when-brighter trend and achromatic phase profile. This favors the geometric interpretation, where the achromatic nature of the global QPO suggests it arises from large-scale geometric changes affecting a steady emission envelope. In contrast, the phase-locked hardening represents localized microphysical processes. This separation of scales implies a two-component jet structure: a soft-spectrum envelope modulated by geometry and a stochastic core where intermittent hard shocks occur.
@@ 5 Summary and Conclusions
By combining the spectral search of Dinesh et al. (2025) with the analysis of the year-scale periodic modulation of PKS 2155-304 over 18 years of observations, we identify a physical connection between jet geometry and transient particle acceleration. Our main conclusions are as follows: (1) PKS 2155-304 shows a softer-when-brighter behavior supporting a geometric origin for its 1.7 yr oscillation. (2) The exceptional spectral hardening event reported in Dinesh et al. (2025) is phase-locked to the QPO trough, coinciding with the minimum of the flux modulation. (3) This alignment supports a Geometric Masking scenario, where the suppression of boosted soft emission at the flux minimum enables the detection of otherwise masked acceleration signatures.

### v2

@@ Abstract
Methods. We analyze first 17.4 years of Fermi-LAT data using 30-day binning. To isolate variability on QPO timescales, we apply Singular Spectrum Analysis to mitigate red-noise effects and use a Moving Block Bootstrap approach to quantify correlations between photon flux and photon index while accounting for temporal autocorrelation.
@@ 1 Introduction
In this Letter, we investigate the physical link between the 1.7 yr QPO and the transient spectral hardening in PKS 2155-304, defining spectral hardening as a departure from a single power-law gamma-ray spectrum, manifested as high-energy upturns in the GeV band. Using 17.4 years of Fermi-LAT data, we characterize the flux vs. photon index behavior and map the hardening event onto the phase of the oscillation. We find that the bulk variability is atypical for an HSP, showing a softer-when-brighter trend with no significant correlation between photon index and phase, consistent with the coexistence of stochastic plasma-driven activity and a geometric modulation. In addition, the uncommon spectral hardening event is phase-locked to the QPO trough.
@@ 2 Data and Methodology
We analyzed the $\gamma$-ray emission of PKS 2155-304 using approximately 17.4 years of Fermi-LAT data (August 2008–January 2026; MJD 54698–61058) obtained from the Fermi-LAT Light Curve Repository [^1] https://fermi.gsfc.nasa.gov/ssc/data/access/lat/LightCurveRepository/ (LCR; Abdollahi et al., 2023). We utilized the standard LCR light curves integrated strictly over the 100 MeV to 100 GeV energy range. As a well-established TeV blazar with a characteristically hard Fermi-LAT spectrum (average photon index $\Gamma\approx 1.8$), the observed $\gamma$-ray flux of PKS 2155-304 is driven by high-energy photons (e.g., Aharonian et al., 2007; Ackermann et al., 2016). To characterize variability on the timescales of the reported QPO, we adopted a 30-day binning with a minimum test statistic threshold of $TS\geq 4$ per bin and allowed the photon index ($\Gamma$) to vary freely in each bin.
@@ 3.1 Global Spectral Behavior
We first characterized the global spectral variability of PKS 2155-304 over the full 17.4-year baseline to establish a physical reference for the QPO modulation (Figure 1). Figure 2 illustrates the relationship between the photon flux and the photon index. In contrast to the harder-when-brighter trend typical of most HSP blazars, PKS 2155-304 shows a statistically significant positive correlation.
@@
Figure 2: Spectral variability of PKS 2155-304. Left: Photon index vs. photon flux, showing a positive correlation ($\rho\approx 0.3$), weak but robust, indicative of an atypical softer-when-brighter trend. Right: Phase-resolved spectral evolution folded at the 1.7 yr period, with the 17.4-year baseline variability (green dots) consistent with an achromatic trend given the lack of significant correlation between photon index and phase ($\rho=0.036$, $p=0.60$), and the two consecutive 30-day bins of the extreme spectral hardening event reported by Dinesh et al. (2025, red dots).
@@ 3.2 Phase-Locked Spectral Hardening
To quantify the significance of this alignment, we performed a phase-scrambling Monte Carlo test ($10^{4}$ realizations) in which the MJD window of the hardening event was randomly reassigned to any valid 30-day bin within the 17.4-year baseline, and the resulting phase was recorded. The probability of a random placement falling within $|\Delta\phi|\leq 0.03$ of the trough center is $p\approx 0.059$. Combining this with the independent flux–index correlation result ($p\approx 1.8\times 10^{-5}$) under the assumption of statistical independence, the joint probability of both features arising by chance is $p_{\mathrm{joint}}\approx 1.1\times 10^{-6}$, strongly disfavoring a coincidental alignment. We interpret the physical origin of this phase-locked behavior in the following section. Note that the assumption of independence is reasonable here, as the global flux-index correlation reflects the overall chromatic behavior of the source across the full 17.4-year dataset, while the phase alignment test specifically evaluates the timing of a single transient hardening event relative to the periodic cycle. These probe distinct physical and statistical aspects of the variability, with no evident coupling in the underlying data generation process.
@@ 4.1 The Geometric Masking Scenario
The phase alignment of the spectral hardening event with the QPO trough ($\phi\approx 0.53$) suggests a conditional visibility mechanism. This trend is naturally explained by geometric Doppler-modulation models, a foundational framework in blazar physics (e.g., Blandford & Rees, 1978; Urry & Padovani, 1995; Rieger, 2004). Relativistic beaming amplifies flux according to $F\propto\delta^{3+\alpha}$, where $\alpha$ is the spectral index ($F_{\nu}\propto\nu^{-\alpha}$) (e.g., Lind & Blandford, 1985; Dermer, 1995). Because the soft emission envelope (e.g., a cooled electron population or jet sheath) has a steeper index than the stochastic hard core, it experiences a stronger differential Doppler boost (e.g., Ghisellini et al., 2005). Consequently, at the flux maximum (smallest viewing angle), the geometrically boosted soft emission overwhelms the hard core, naturally producing the observed softer-when-brighter trend (e.g., Villata et al., 2004; Raiteri et al., 2017). The presence of such active Doppler factor variations in PKS 2155-304 has been independently confirmed by contemporary multi-epoch SED modeling (Harutyunyan et al., 2026). We emphasize that this is a phenomenological application of the same geometric framework recently utilized to interpret the QPO in PG 1553+113 (Madero & Domínguez, 2026).
In this Geometric Masking scenario, we propose that the observational rarity of spectral hardening in HSP blazars ($0.1\%$; Dinesh et al. 2025) is not necessarily an indication that extreme particle acceleration events are uncommon, but rather that they are frequently masked. In a geometric QPO scenario, flux peaks are driven by maximum relativistic beaming. Our finding of a global softer-when-brighter trend ($\rho\approx 0.3$) indicates that these boosted states are dominated by a soft photon field, likely originating from cooled electron populations or the jet sheath. This may redefine the high-flux state not as a period of intrinsic quiescence, but as a regime of maximum geometric masking. The jet may host frequent plasma-driven particle acceleration events that remain observationally unresolved, overwhelmed by the intensified soft background. During these phases, the spectral signature of intrinsic acceleration, such as shocks or magnetic reconnection, is suppressed by the boosted soft emission. Conversely, at the QPO trough, the Doppler factor is minimized, effectively suppressing the geometric boost and reducing this soft background. This reduces the geometric masking, allowing hard-spectrum emission from transient jet instabilities to dominate the observed flux and making the hardening detectable.
@@
An important corollary of this scenario addresses the rarity of the detected event. While the masking effect is minimized periodically every 1.7 yr, Dinesh et al. (2025) reported only one significant spectral hardening episode over the 17.4-year baseline. This apparent scarcity is likely due to two factors. First, the geometric minimum is a necessary condition for visibility, not a driver of acceleration; an intrinsic stochastic shock must still coincide with the narrow temporal window of the trough. Second, the blind search strategy employed by Dinesh et al. (2025) relied on high-significance deviations which favor high-flux states. Since QPO troughs are by definition low-flux intervals, weaker or shorter-duration hardening events may have failed to meet the strict global detection thresholds. We predict that a targeted search specifically focused on the QPO troughs of PKS 2155-304 will reveal a population of sub-threshold hardening events that were missed by previous blind surveys.
@@ 4.2 Constraints on Jet Structure and Extreme States
The coexistence of a stable geometric modulation and transient spectral hardening provides a unique constraint on the jet’s spatial architecture. While purely plasma-driven QPO models, such as recurrent shocks (e.g., Kirk et al., 1998), could in principle produce similar periodicities, they generally predict harder-when-brighter spectral evolution, which is inconsistent with our observed softer-when-brighter trend and achromatic phase profile. This favors the geometric interpretation, where the achromatic nature of the global QPO suggests it arises from large-scale geometric changes affecting a steady emission envelope. In contrast, the phase-locked hardening represents localized microphysical processes. This separation of scales implies a two-component jet structure: a soft-spectrum envelope modulated by geometry and a stochastic core where intermittent hard shocks occur. Recent time-resolved SED modeling of PKS 2155-304 over the same epoch supports this duality, revealing that while many flaring states are driven by bulk Doppler factor variations, isolated gamma-ray flares require localized hardening of the injected electron distribution without bulk geometric changes (Harutyunyan et al., 2026).
@@
4.3 Comparison with accretion-shock resonance models
Alternative models, such as accretion-shock resonance oscillations (e.g., Molteni et al., 1996; Dash et al., 2026), can produce phase-locked spectral changes via coronal thermal Comptonization in radio-quiet active galactic nuclei and X-ray binaries. However, PKS 2155-304 is a highly beamed blazar whose >100>100 MeV gamma-ray emission is generally attributed to non-thermal Synchrotron Self-Compton (SSC) processes within the relativistic jet (e.g., Costamante & Ghisellini, 2002; Katarzyński et al., 2008; Kusunose & Takahara, 2008; Aharonian et al., 2009). Recent broadband SED modeling over the 2008–2023 timeframe finds that the emission is well-described by a leptonic SSC jet model, making accretion disk thermal Comptonization a less likely primary driver for the gamma-ray flux (Harutyunyan et al., 2026). Furthermore, accretion-shock models typically predict a continuous, systematic phase-dependent spectral evolution (e.g., persistently hardest at phase $\sim 0.5$). The overall achromatic QPO phase profile in our data ($\rho=0.036$), punctuated by a single rare hardening event, is less consistent with a persistent disk-shock driver and aligns more closely with geometric Doppler variations.
@@ 5 Summary and Conclusions
By combining the spectral search of Dinesh et al. (2025) with the analysis of the year-scale periodic modulation of PKS 2155-304 over 17.4 years of observations, we identify a physical connection between jet geometry and transient particle acceleration. Our main conclusions are as follows: (1) PKS 2155-304 shows a softer-when-brighter behavior supporting a geometric origin for its 1.7 yr oscillation. (2) The exceptional spectral hardening event reported in Dinesh et al. (2025) is phase-locked to the QPO trough, coinciding with the minimum of the flux modulation. (3) This alignment supports a Geometric Masking scenario, where the suppression of boosted soft emission at the flux minimum enables the detection of otherwise masked acceleration signatures.
@@
Acknowledgements.
This work has made use of public Fermi data obtained from the High Energy Astrophysics Science Archive Research Center (HEASARC), provided by NASA Goddard Space Flight Center. We specifically acknowledge the Fermi Science Support Center (FSSC) for the provision of the Fermi-LAT Light Curve Repository.